OpenStudy (anonymous):

2[(2y−1)+y]≤5(y+3) can someone help me solve this

OpenStudy (dumbcow):

distribute first of all \[6y -2 \le 5y +15\] then combine like terms 6y -5y = y 15+2 = 17 \[y \le 17\]

OpenStudy (anonymous):

idk the order to distribte in original equation like what side to start on

OpenStudy (anonymous):

well right side is easy left side i dont understand

OpenStudy (dumbcow):

ignore the parenthesis around "2y-1" ---> 2( 2y -1 +y) = 4y +2 +2y = 6y + 2

OpenStudy (anonymous):

ok thx easy now